OpenStudy (anonymous):

Find the area of a circle whose diameter is 4.2 km. Round to the nearest hundredths. A. 6.59 square km B. 13.19 square km C. 13.85 square km D. 55.39 square km

OpenStudy (anonymous):

answer please

OpenStudy (texaschic101):

A = pi * r^2 r = diameter/2 = 4.2/2 = 2.1 A = 3.14 * 2.1^2 A = 3.14 * 4.41 A = 13.8474 rounded would be 13.85